A person walks70 m east,then 30 m west.
what is his distance travelled?
what is his displacement?
if the trip lasted 20 seconds,
what was his velocity?
what was his average speed?

Answer by Alan3354(69443)

A person walks 70 m east, then 30 m west.
what is his distance travelled?
70 + 30 = 100 meters
-----
what is his displacement?
70 - 30 = 40 meters
-----
if the trip lasted 20 seconds,
what was his velocity?
Assuming constant speed:
For 14 seconds, it was 5 m/sec East
For 6 seconds, it was 5 m/sec West
----------
what was his average speed?
100/20 = 5 m/sec